You are given the production function: 𝑄(𝐾, 𝐿) = 10𝐾 𝛼𝐿 𝛽 a) What is the average product of labour, holding capital fixed at K? Simplify fully [6 marks] b) What is the marginal rate of technical substitution (MRTS)? Simplify fully [13 marks] Does the above function exhibit increasing, decreasing or constant returns to scale? Illustrate why and explain what this means [6 marks


1
Expert's answer
2021-05-06T13:38:04-0400

Average product of labor (APL), is total product (TP) divided by the total quantity of labor.

Given;

production function

"Q(K,L)=10K^{\alpha}L^{\beta}"

(a)

Average product of labor holding capital fixed at K.

"=\frac{Production(Q)}{Labor(L)}"


"=\frac{10K^{\alpha}L^{\beta}}{L}"


Average Product of Labor (APL)"=10K^{\alpha}L^{\beta -1}"


(b)

Marginal rate of technical substitution illustrates the rate at which one factor must decrease so that the same level of productivity can be maintained when another factor is increased.

"MRTS {K \atop L} =\frac{MP_L}{MP_k}=\frac{\delta Q \delta L}{\delta Q \delta K}=\frac{10\beta K^{\alpha}L^{\beta -1}}{10\alpha K^{\alpha -1}L^{L^{\beta}}}"


"MRTS=\frac{\beta}{\alpha}\frac{K}{L}"


As, production "f^n:Q=10K^{\alpha}L^{\beta}."

"{let \lambda >1}"

"Q(\lambda K, \lambda L)=10(\lambda K)^{\alpha}(\lambda L)^{\beta}"

"=10\lambda ^{\alpha +\beta}(K^{\alpha L^{\beta}})"

"[Q(\lambda K, \lambda L) =\lambda ^{\alpha +\beta}Q(K,L).]"


If "\alpha + \beta=1," the function exhibits constant returns to scale.

If "\alpha + \beta >1," the function exhibits increasing returns to scale.

if "\alpha + \beta<1," the function exhibits decreasing returns to scale.


It means:

1. If inputs are increased by "\lambda" (some) amount, then the output increases "\lambda" amount., then it is called constant returns to scale.

2. If inputs are increased by "\lambda" (some) amount, then the output increases by more than "\lambda" amount., then it is called increasing returns to scale.

3. If inputs are increased by "\lambda" (some) amount, then the output increases by less than "\lambda" amount., then it is called decreasing returns to scale.